Can You Help Me???
Please Simplify:
3x^2-22x+7
: Can You Help Me???
Please Simplify:
3x^2-22x+7

Answer by Fombitz(1756)

Instead of simplify, I think you mean factor.
3x^2-22x+7=(3x+a)(x+b)
Where a*b=7 and a+3b=-22.
Let's try a=-1, b=-7.
That satisfies both conditions.
3x^2-22x+7=(3x-1)(x-7)
Question 149362: Can You Help Me???
Please Simplify:
3x^2-22x+7
: Can You Help Me???
Please Simplify:
3x^2-22x+7

Answer by ankor@dixie-net.com(4538)

Please Simplify:
3x^2-22x+7
:
You can factor this to:
(3x - 1)(x - 7)